OpenStudy (telly):

the jogging track is 5/9 of a mile long. If Ashley jogged around it 4 times, how far did she run?

OpenStudy (anonymous):

\[\frac{ 5 }{ 9 } \times \frac{ 4 }{ 1 } = \frac{ 20 }{ 9}= 2\frac{ 2 }{ 9 }\quad or \quad 2.222 \quad miles\]

OpenStudy (anonymous):

And Welcome to OpenStudy ^.^